OpenStudy (anonymous):

Solve by using the perfect squares method. x2 + 16x + 64= 0

OpenStudy (anonymous):

@mathmate

pooja195 (pooja195):

\(\huge\color{orange}{Welcome~To~OpenStudy}\) Hey there!!! Since you are new here, read this legendary tutorial for new OpenStudiers!! http://goo.gl/5pp1u0 For an visual tutorial look at this! Its amazing :o https://prezi.com/fs3hqdpcopic/an-unofficial-guide-to-openstudy/ Also consider reading the CoC and the T&C http://openstudy.com/code-of-conduct http://openstudy.com/terms-and-conditions

pooja195 (pooja195):

\[\huge x^2 + \sqrt{16x}+ \sqrt{64}= 0\]

OpenStudy (anonymous):

what is that

pooja195 (pooja195):

we square root the perfect squares

OpenStudy (mathmate):

@pooja195 We squareroot the first and last terms. \(\sqrt{x^2}=x\) and \(\sqrt{64}=8\)

OpenStudy (mathmate):

So far so good, the first and last terms are perfect squares. Next:

OpenStudy (mathmate):

we double the product, 2(8x)=16x If this equals the absolute value of the middle term |16x|, we have a perfect square expression. The sign between x and 8 is determined by the sign of the middle term.

OpenStudy (mathmate):

So \(x^2 + 16x + 64= (x+8)^2\), or \((x+8)^2=0\)

OpenStudy (mathmate):

@Nitaoffaith I'll let you take it from here to finish solving the equation.

OpenStudy (anonymous):

ok thx

OpenStudy (mathmate):

You're welcome! :)